# figureb tableb grand_reference.bib

### v1

@@ V-A1 Step-wise optimizations for K-means
Figure 7 demonstrates how our K-Means distance kernel is optimized from 5% to 182% of cuML stepwise. The performance is measured with GFLOPS (bar plot, the left y-axis) and the performance ratio with respect to cuML (line chart, the right y-axis). Without using any optimizations, the K-Means Naive obtains a performance of 482 GFLOPS. Next K-Means V1 employs GEMM to K-Means distance calculation. The performance improved from 482 GFLOPS to 4662 GFLOPS. And then, K-Means V2 adds kernel fusion to both thread and threadblock level, reducing memory operations. The performance is improved to 5902 GFLOPS. After that, we apply threadblock level broadcast to further reduce memory bound. The performance of K-Means V3 achieves 6916 GFLOPS. And tensor core enabled and parameter selection, we finally acheive 17686 GFLOPS, which exceeded the performance of cuML (9676 GFLOPS).
@@ V-A2 Performance evaluation with M and K fixed
Figure 8 and 9 demonstrate a performance evaluation in distance step between FT K-Means, two selected parameters relative to cuML (all without fault tolerance) in FP32 and FP64 precision when M and K are fixed. We tested the performance of different methods under two values of K, $K=8$ and $K=128$. These are cases representing two situations: one with very few dimensions of data and the other one with relatively more dimensions of data. The selected parameters are labeled Parameter1 and Parameter2, and they are chosen based on experience. And the same parameter name refers to different parameters in FP32 and FP64, but their performance is similar. For both FP32 and FP64 precision, parameters selected through experience cannot achieve good performance. Parameter 1 is always slower than cuML, with an average overhead of 15%. And for parameter 2, it slightly exceeded the performance of cuML when $K=8$ in FP32, and it achieves the performance of cuML in some data points where N is small. However, its overall performance is still 5% slower than cuML. Using code generation strategy, we achieved 235% speedup compared to cuML under FP32 precision, and the gain in performance is significant even when K is relatively larger. However, under FP64 precision, improvements in performance is minimal, with a overall speedup of 4% in these two cases. And the curves of cuML and FT K-Means are almost coincident in a large portion of data points.
Figure 10: FP32 precision comparison of K-Means performance at distance step without fault tolerance with FT K-Means, Selected parameters and cuML on an A100 GPU, with M and N fixed
Figure 11: FP64 precision comparison of K-Means performance at distance step without fault tolerance with FT K-Means, Selected parameters and cuML on an A100 GPU, with M and N fixed
@@ V-A3 Performance evaluation with M and N fixed
Figure 10 and 11 offer a performance evaluation in distance step between FT K-Means, two selected parameters relative to cuML (all without fault tolerance) in FP32 and FP64 precision when M and N are fixed. The main parameter settings are similar to the previous section, and $N=8$ and $N=128$ indicate fewer clustering centroids and relatively more clustering centroids. For both FP32 and FP64 precision, parameters selected through experience cannot achieve good performance. Parameter 1 has an average overhead of 30%. And parameter 2 exceeded the performance of cuML when K is small in some FP32 cases, and outperforms cuML in FP64 when $N=8$ (1.03x speedup). but its overall the performance is still 15% slower than cuML. Using code generation strategy, we obtained 239% speedup compared to cuML under FP32 precision, which is similar to fixing M and K. Under FP64 precision, improvements in performance is higher than previous section, which is 8% in these two cases. And the speedup in small N is relatively considerable (15%).
@@ V-A4 Overall performance evaluation
Figure 12 illustrate an overall comparison between our code generation method and cuML K-means on both FP32 and FP64. For FP32. Our approach shows significant improvement in performance, with a maximum speedup of 4.55x and an average speedup of 2.49x. From the perspective of dimension N, the figure indicates a clear trend that the performance improvement diminishes. Furthermore, $N=64$ serves as a clear threshold, beyond which speedup essentially decreases to below 2.0x. Such trend becomes even more manifest in the case of FP64. The average speedup is only 1.04x, with a maximum speedup of 1.39x. When N exceeds 32, the performance of our method drops to almost identical as cuML. Meanwhile, there is no apparent trend observed in dimension K for both FP32 and FP64.
@@ V-A5 Evaluation of parameter selection
The experimental results from previous sections motivate us to analyze the parameters chosen for our code generation method for each data point. For the data size of Figure 12, we generated 120 groups of FP32 parameters and 80 groups of FP64 parameters to be selected from. However, only 7 groups of FP32 parameters and 4 groups of FP64 parameters are actually chosen in at least one data point. And all threadblock and warp level parameters are shown in Figure 13. And thread level parameters are fixed (FP32: 16,8,4 FP64: 8,8,4) owing to the constant size of tensor core. Moreover, we depicted figure 14 to illustrate the relationship between each data point and its corresponding selected parameters. We will analyze in detail the relationship between the optimal parameters generated by the code generator and the cuML parameters in next section.
@@
V-A6 Detailed analyse of parameters
For FP32, as we can see from figure 14, the data points can be divided into three parts: $N\leq 32$, $32<N\leq 64$ and $64<N$. When $N\leq 32$, one of the optimal parameters is parameter 88. As table I indicates, compared to cuML, it has a bigger threadblock and warp size on M direction and smaller size on N direction. And that is reasonable due to small N size in these data points. threadblock.N = 256 is too big that more than 87.5% of the threadblock size is containing blank data, so the occupancy is very low. Furthermore, warp.N of cuML, which is 64, exceeds the size of N. Therefore the improvement of our new parameters is significant. As N increases, i.e. $32<N\leq 64$, the warp size of cuML becomes justified, and our parameter (parameter 69) sticks to this size. However, the threadblock size remains too big on N direction, so a more balanced version of threadblock $64,128,16$ outperforms. When N keeps increasing, although the occcupancy of cuML achieves a reasonable scope, a more balanced threadblock size, e.g parameter 83 can reduce the overall amount of data movement form GPU memory to shared memory, and therefore improve performance.
For FP64, the data points only have two main parts: $N\leq 32$ and $32<N$. When N is relatively small, parameter21 has advantage in its small Threadblock.N, and increases occupancy. And as N increases, the parameters which are balanced in M and N direction has better performance. And as shown in table I, our parameter 19 is actually identical to cuML’s parameter, which demonstrates that the best parameter choice is cuML’s parameter in these cases. This illustrates an interesting phenomenon: the parameter choices in FP32 are much more than which in FP64. Moreover, there is also greater potential for performance improvement in FP32 than FP64. And there are several reasons for this.
@@ V-A6 Detailed analyse of parameters
First, CUTLASS enables TF32 in the FP32 GEMM kernel, which improves the processing speed of tensor cores. Therefore, the overhead of data movement and epilogue (row-wise reduction) becomes more critical, resulting in greater potential for alternative parameters. Second, the memory alignment requirement for FP64 is more strict than FP32 and is fixed to 1 in CUTLASS’s implementation. So the degree of vectorization for FP64 is lower. So a balanced data fetching pattern (Threadblock.M = Threadblock.N) is crucial for increasing performance. Third, the thread level parameters of FP64 is smaller than which of FP32. Therefore even when N is relatively small, it is easier for a general parameter to obtain high occupancy.
@@ V-C FT K-Means under Error Injections
As illustrated in Figure 17, the experimental evaluation of the FT K-Means algorithm under error injection reveals a minimal average overhead of approximately 2.36%, demonstrating the algorithm’s efficiency in handling errors with minimal additional computational cost. The results indicate robust performance across various input shapes, with overhead percentages ranging from a slight reduction of about -0.93% to a modest increase up to 9.49%. This low overhead highlights the effectiveness of the fault tolerance mechanisms integrated into the FT-KMeans, ensuring correctness even in the presence of injected errors. Kosaian’s FT scheme results in a 100 % overhead due to time redundant recomputation. Wu’s FT scheme introduces an overhead of 30% due to a suboptimal GEMM baseline without using the asynchronous memory copy.
@@ V-D Performance Evaluation on T4
Figure 19 demonstrate a performance evaluation in distance step between FT K-Means, two selected parameters relative to cuML (all without fault tolerance) in FP32 precision when M and K are fixed. The definitions of Parameter1 and Parameter2 are similar to our evaluation on A100. And they are consistent with the values on A100 to the greatest extent. However, they have better performance compared to cuML’s parameter in this architecture, with speed up of 184% and 208%. Using code generation strategy, we achieved 413% speedup compared to cuML under FP32 precision, and the gain in performance is significant even when K is relatively larger.
Figure 19: FP32 precision comparison of K-Means performance at distance step without fault tolerance with FT K-Means, Selected parameters and cuML on a T4 GPU, with M and K fixed.
Figure 20 offers a performance evaluation in distance step between FT K-Means, two selected parameters relative to cuML (all without fault tolerance) in FP32 precision when M and N are fixed. The main parameter settings are similar to the previous section, and $N=8$ and $N=128$ indicate fewer clustering centroids and relatively more clustering centroids. The selected parameters acheive 183% and 206% speed up against cuML. Using code generation strategy, we obtained 381% speedup compared to cuML under FP32 precision, which is similar to fixing M and K.
Figure 20: FP32 precision comparison of K-Means performance at distance step without fault tolerance with FT K-Means, Selected parameters and cuML on a T4 GPU, with M and N fixed
Figure 21 benchmarks the performance of FT K-Means with or without fault tolerance for single precision. FT K-Means shows an average overhead of $18\%$ with fault tolerance and $30\%$ under error injection. Compared with Kosaian’s ABFT, FT K-means has a $60\%$ improvement compared to the combination of Kosaian’s ABFT and time-redundant recomputation. FT K-Means shows a similar performance compared to Wu’s ABFT scheme.
@@ Acknowledgement
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through the Advanced Computing (SciDAC) program under Award Number DE-SC0022209.

### v2

@@ V-A1 Step-wise optimizations for K-means
Figure 7 demonstrates how our K-means distance kernel is optimized from 5% to 182% of cuML stepwise. The performance is measured with GFLOPS (bar plot, the left y-axis) and the performance ratio with respect to cuML (line chart, the right y-axis). Without using any optimizations, the K-means Naive obtains a performance of 482 GFLOPS. Next K-means V1 employs GEMM to K-means distance calculation. The performance improved from 482 GFLOPS to 4662 GFLOPS. Then, K-means V2 adds kernel fusion to both thread and threadblock levels, reducing memory operations. The performance is improved to 5902 GFLOPS. After that, we apply threadblock level broadcast to further reduce memory bound. The performance of K-means V3 achieves 6916 GFLOPS. With tensor core enabled and parameter selection, we finally achieved 17686 GFLOPS, which exceeded the performance of cuML (9676 GFLOPS).
@@ V-A2 Performance evaluation with M and K fixed
Figure 8 and 9 demonstrate a performance evaluation in distance step between FT K-means, two selected parameters relative to cuML (all without fault tolerance) in FP32 and FP64 precision when M and K are fixed. We tested the performance of different methods under two values of K, $K=8$ and $K=128$. These are cases representing two situations: one with very few dimensions of data and the other one with relatively more dimensions of data. The selected parameters are labeled Parameter1 and Parameter2, and they are chosen based on experience. The same parameter name refers to different parameters in FP32 and FP64, but their performance is similar. For both FP32 and FP64 precision, parameters selected through experience cannot achieve good performance. Parameter 1 is always slower than cuML, with an average overhead of 15%. For parameter 2, it slightly exceeded the performance of cuML when $K=8$ in FP32, and it achieves the performance of cuML in some data points where N is small. However, its overall performance is still 5% slower than cuML. Using code generation strategy, we achieved 235% speedup compared to cuML under FP32 precision, and the gain in performance is significant even when K is relatively larger. However, under FP64 precision, improvements in performance are minimal, with an overall speedup of 4% in these two cases. The curves of cuML and FT K-means are almost coincident in a large portion of data points.
Figure 10: FP32 precision comparison of K-means performance at distance step without fault tolerance with FT K-means, selected parameters, and cuML on an A100 GPU, with M and N fixed
Figure 11: FP64 precision comparison of K-means performance at distance step without fault tolerance with FT K-means, Selected parameters, and cuML on an A100 GPU, with M and N fixed
@@ V-A3 Performance evaluation with M and N fixed
Figure 10 and 11 offer a performance evaluation in distance step between FT K-means, two selected parameters relative to cuML (all without fault tolerance) in FP32 and FP64 precision when M and N are fixed. The main parameter settings are similar to the previous section, and $N=8$ and $N=128$ indicate fewer clustering centroids and relatively more clustering centroids. For both FP32 and FP64 precision, parameters selected through experience cannot achieve good performance. Parameter 1 has an average overhead of 30%. Parameter 2 exceeded the performance of cuML when K is small in some FP32 cases, and outperforms cuML in FP64 when $N=8$ (1.03x speedup). but its overall performance is still 15% slower than cuML. Using code generation strategy, we obtained 239% speedup compared to cuML under FP32 precision, which is similar to fixing M and K. Under FP64 precision, performance improvements are higher than in the previous section, which is 8% in these two cases. And the speedup in small N is relatively considerable (15%).
@@ V-A4 Overall performance evaluation
Figure 12 illustrates an overall comparison between our code generation method and cuML K-means on both FP32 and FP64. For FP32. Our approach shows significant performance improvement, with a maximum speedup of 4.55x and an average speedup of 2.49x. From the perspective of dimension N, the figure indicates a clear trend that the performance improvement diminishes. Furthermore, $N=64$ serves as a clear threshold, beyond which speedup essentially decreases to below 2.0x. The marginal improvement becomes even more common in the case of FP64. The average speedup is only 1.04x, with a maximum speedup of 1.39x. When N exceeds 32, the performance of our method drops to almost identical to cuML. Meanwhile, there is no apparent trend observed in dimension K for both FP32 and FP64.
@@ V-A5 Evaluation of parameter selection
The experimental results from previous sections motivate us to analyze the parameters chosen for our code generation method for each data point. For the data size of Figure 12, we generated 120 groups of FP32 parameters and 80 groups of FP64 parameters to be selected from. However, only 7 groups of FP32 parameters and 4 groups of FP64 parameters are actually chosen in at least one data point. All threadblock and warp level parameters are shown in Figure 13. And thread level parameters are fixed (FP32: 16,8,4 FP64: 8,8,4) owing to the limited size of the tensor core. Moreover, figure 14 illustrates the relationship between each data point and its corresponding selected parameters. We will analyze in detail the relationship between the optimal parameters generated by the code generator and the cuML parameters in the next section.
@@
V-A6 Detailed analysis of parameters
For FP32, as shown in figure 14, the data points can be divided into three parts: $N\leq 32$, $32<N\leq 64$ and $64<N$. When $N\leq 32$, one of the optimal parameters is parameter 88. As table I indicates, compared to cuML, it has a bigger threadblock and warp size in the M direction and a smaller size in the N direction. That is reasonable due to the small N size in these data points. threadblock. N = 256 is too big that more than 87.5% of the threadblock size contains blank data, so the occupancy is very low. Furthermore, warp.N of cuML, which is 64, exceeds the size of N. Therefore the improvement of our new parameters is significant. As N increases, i.e. $32<N\leq 64$, the warp size of cuML becomes justified, and our parameter (parameter 69) sticks to this size. However, the threadblock size remains too big in N direction, so a more balanced version of threadblock $64,128,16$ outperforms. When N keeps increasing, although the occupancy of cuML achieves a reasonable scope, a more balanced threadblock size, e.g. parameter 83 can reduce the overall amount of data movement from GPU memory to shared memory, and therefore improve performance.
For FP64, the data points only have two main parts: $N\leq 32$ and $32<N$. When N is relatively small, parameter 21 has an advantage in its small Threadblock. N, and increases occupancy. As N increases, the parameters that are balanced in the M and N direction have better performance. As shown in table I, our parameter 19 is actually identical to cuML’s parameter, which demonstrates that the best parameter choice is cuML’s parameter in these cases. This illustrates an interesting phenomenon: the parameter choices in FP32 are much more than those in FP64. Moreover, there is also greater potential for performance improvement in FP32 than in FP64. There are several reasons for this.
@@ V-A6 Detailed analysis of parameters
First, CUTLASS enables TF32 in the FP32 GEMM kernel, which improves the processing speed of tensor cores. Therefore, the overhead of data movement and epilogue (row-wise reduction) becomes more critical, resulting in greater potential for alternative parameters. Second, the memory alignment requirement for FP64 is more strict than FP32 and is fixed to 1 in CUTLASS’s implementation. So the degree of vectorization for FP64 is lower. So a balanced data fetching pattern (Threadblock.M = Threadblock.N) is crucial for increasing performance. Third, the thread-level parameters of FP64 are smaller than those of FP32. Therefore even when N is relatively small, it is easier for a general parameter to obtain high occupancy. In the future, deep reinforcement learning can help accelerate the hyperparameter tuning [johnston2023curriculum, liu2023stationary, johnston2023downlink].
@@ V-C FT K-means under Error Injections
As illustrated in Figure 17, the experimental evaluation of the FT K-means algorithm under error injection reveals a minimal average overhead of approximately 2.36%, demonstrating the algorithm’s efficiency in handling errors with minimal additional computational cost. The results indicate robust performance across various input shapes, with overhead percentages ranging from a slight reduction of about -0.93% to a modest increase up to 9.49%. This low overhead highlights the effectiveness of the fault tolerance mechanisms integrated into the FT-KMeans, ensuring correctness even in the presence of injected errors. Wu’s FT scheme introduces an overhead of 30% due to a suboptimal GEMM baseline without using the asynchronous memory copy.
@@ V-D Performance Evaluation on T4
Figure 19 demonstrates a performance evaluation in distance step between FT K-means, two selected parameters relative to cuML (all without fault tolerance) in FP32 precision when M and K are fixed. The definitions of Parameter1 and Parameter2 are similar to our evaluation of A100. And they are consistent with the values on A100 to the greatest extent. However, they have better performance compared to cuML’s parameter in this architecture, with speeds up to 184% and 208%. Using code generation strategy, we achieved 413% speedup compared to cuML under FP32 precision, and the gain in performance is significant even when K is relatively larger.
Figure 19: FP32 precision comparison of K-means performance at distance step without fault tolerance with FT K-means, Selected parameters, and cuML on a T4 GPU, with M and K fixed.
Figure 20 offers a performance evaluation in distance step between FT K-means, two selected parameters relative to cuML (all without fault tolerance) in FP32 precision when M and N are fixed. The main parameter settings are similar to the previous section, and $N=8$ and $N=128$ indicate fewer clustering centroids and relatively more clustering centroids. The selected parameters achieve 183% and 206% speed up against cuML. Using code generation strategy, we obtained 381% speedup compared to cuML under FP32 precision, which is similar to fixing M and K.
Figure 20: FP32 precision comparison of K-means performance at distance step without fault tolerance with FT K-means, Selected parameters, and cuML on a T4 GPU, with M and N fixed
Figure 21 benchmarks the performance of FT K-means with or without fault tolerance for single precision. FT K-means shows an average overhead of $18\%$ with fault tolerance and $30\%$ under error injection. Compared with Wu’s ABFT, FT K-means has a $60\%$ improvement due to the elimination of threadblock-level synchronization.
@@ Acknowledgement
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, and Scientific Discovery through the Advanced Computing (SciDAC) program under Award Number DE-SC0022209.